# KIX: A Knowledge and Interaction-Centric Metacognitive Framework for Task Generalization

### v2

@@ Abstract
People aptly exhibit general intelligence behaviors in solving a variety of tasks with flexibility and ability to adapt to novel situations by reusing and applying high-level knowledge acquired over time. But artificial agents are more like specialists, lacking such generalist behaviors. Artificial agents will require understanding and exploiting critical structured knowledge representations. We present a metacognitive generalization framework, Knowledge-Interaction-eXecution (KIX), and argue that interactions with objects leveraging type space facilitate the learning of transferable interaction concepts and generalization. It is a natural way of integrating knowledge into reinforcement learning and is promising to act as an enabler for autonomous and generalist behaviors in artificial intelligence systems.
1 Introduction
Figure 1: Learning generalist behaviors requires developing a natural and abstract understanding of the environment, as well as meta-reasoning and interaction abilities. The metacognitive knowledge-guided approach exploits inherent abstractions and relationships between objects to generate interaction goals, enabling the agent to interact with objects and generalize to solve complex tasks.
@@ 1 Introduction
Flexibility, ability to actively think and evaluate options –learning and reasoning to enable adaptation– are hallmarks of human intelligence. Core to this intelligence is how easily people can find natural decomposition in world environments and tasks; this enables flexible, generalist behaviors. Replicating such autonomy and general intelligence abilities in artificial agents continues to pose significant challenges, which we argue requires a knowledge and interaction-centric metacognitive approach.
@@
Meta-reasoning about next actions, their values, and possible next states facilitates flexible decision-making in novel situations, problem solving, concept learning, and knowledge acquisition. This reasoning extends beyond mere memory lookup. Meta-reasoning is more broadly sense-making in a knowledge system - enabling people to intuitively direct their learning to make sense of relationships between entities, interact with the same entities in diverse ways, and ground their interactions for both short- and long-term goals. Metacognitive reasoning appears to be central to generalization. In contrast, meta-reasoning systems in artificial intelligence are primarily focused on resource allocation problems. While resource allocation underpins efficient functioning, it is crucial to ideate metacognitive generalization, Figure 1, to enable generalist behaviors in artificial agents.
@@
Our proposal builds on three main principles for solving complex tasks flexibly. Natural Decomposition and Abstraction: World environments are naturally composed of entities categorized by their types, similar to a natural hierarchy of objects and classes. Agents interact with entities in the instance space, which inherit properties and relationships from a higher-level type space. Interactions: World environments are essentially knowledge graphs representing underlying objects and their relationships. An agent’s metacognitive process employs these knowledge graphs to create a plan specifying which entities to interact with and the nature of these interactions to complete tasks and achieve goals. Composition of Interactions: Agents’ flexible behaviors stem from their ability to learn interactions and leverage interaction pathways for graph traversal at both the type level and the instance level, enabling quick adaptation. Integrating these principles into metacognitive learning, reasoning, and control provides the levers and connecting points for instilling generalist behaviors in agents.
@@
The proposed metacognitive framework captures the essence of generalist behaviors and may have far-reaching implications across various domains. In this article, we describe the key components of the framework that exemplify generalist characteristics, followed by an experimental evaluation as well as how we envision it enabling flexibility through meta-level knowledge guided interactions.
2 Background
@@
Compositionality and abstractions are central to generalization learning, with abstraction requiring a mechanism for grounding representations in actual objects. In previous work, Veerapaneni et al. (2020) presented an object-centric model where object representations were constructed from raw observations and simpler local functions with uniformity, which were then used for prediction and planning. Similarly, Vijay et al. (2019) used a symbolic object-centric representation to study the impact of adding prior knowledge through relations in an object graph. These studies focus primarily on object space, which inhibits the abstraction of objects and relational structures naturally occurring in the real world and used for decision-making. Key questions include what a natural abstraction should be and how it can be used to compose interactions as building blocks leveraging natural relational structures. We address these largely overlooked questions by preserving abstract knowledge and relational higher-order structures in the form of knowledge graphs, where abstractions are grounded using type space. Learning a compositional structure is achieved through learning over knowledge graphs, with neighborhood aggregation until stability is reached. This distinguishes our approach from previous works.
Other methods emphasize different aspects, such as encouraging exploration through intrinsic reward or information gain Ten et al. (2022); Oudeyer et al. (2007); Pathak et al. (2017), motivation Moerland et al. (2023); Zhu et al. (2023), developing internal goals for empowerment Klyubin et al. (2005), using a curriculum to progress from easier to more difficult tasks Soviany et al. (2022); Campero et al. (2020), or inducing a learning hierarchy based on teacher signals Portelas et al. (2020) Pateria et al. (2021); Hutsebaut-Buysse et al. (2022) or rewards Furelos-Blanco et al. (2023). These methods focus on the execution space, influencing reward signals for learning policies, but largely ignore the higher-order structures and relations between objects in the environment. Our work distinguishes itself by incorporating natural relational structures through knowledge graphs and leverages meta-level knowledge structures to guide interactions that govern low-level execution.
@@
Meta-reasoning is often defined as the set of processes that monitor our thought processes and regulate resource allocation Ackerman and Thompson (2017). Its basic structure Nelson (1990) consists of two levels: meta and object. The meta level’s functions include control and monitoring. The object’s state changes as control passes from the meta to the object level. The monitoring process allows the meta controller to track the status of object-level processes and actions, as well as the consequences of those actions. However, current meta-cognitive reasoning, primarily as a resource manager, is insufficient for exhibiting generalist behaviors, necessitating generalist meta-cognition approaches such as the one proposed here.
@@
$h_{v}^{(k)}=\mathcal{F}^{\prime}\left(h_{v}^{(k-1)},\bigoplus_{u\in\mathcal{N}(v)}\mathcal{F}^{\prime\prime}\left(h_{v}^{(k-1)},h_{u}^{(k-1)},\mathtt{e}_{u,v}\right)\right)$
where $h_{v}^{(k)}$ is the updated representation of node $v$ embedding in layer $k$, $\mathcal{N}(v)$ is the set of neighboring nodes of node $v$, $h_{u}^{(k-1)}$ is the representation of neighboring node $u$ in layer $k-1$, $\bigoplus$ is a differentiable, permutation invariant aggregation function, $\mathcal{F}^{\prime}$ and $\mathcal{F}^{\prime\prime}$ are differentiable functions usually neural networks. Node features are transformed prior to aggregation, and the methods used to transform and aggregate features differ between graph network variants. Motivated by the attention mechanism Vaswani et al. (2017) in the transformer model, Veličković et al. (2017) employs self-attention over node features and aggregating over neighborhood nodes in graph attention networks (GAT). We use the GATv2 operator Brody et al. (2021), which improves static attention problems in graph attention layers.
@@
We consider a standard reinforcement learning setting Sutton and Barto (2018) in which an agent interacts with the environment over a number of steps. A Markov decision process (MDP) is defined as a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma)$ in which $\mathcal{S}$ is a finite state space, $\mathcal{A}$ is a finite discrete action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ is the transition model, $r:\mathcal{S}\times\mathcal{S}\rightarrow\mathbb{R}$ is the reward function and $\gamma\in(0,1)$ is the discount factor. At each step t, the agent takes an action $a_{t}\in\mathcal{A}$ in state $s_{t}\in\mathcal{S}$ according to its policy $\pi$, which is a mapping of states to actions. As a result, the agent reaches state $s_{t+1}\in\mathcal{S}$, receives a scalar reward $r_{t}\in\mathbb{R}$, and accumulates returns; $R_{t}=\sum_{k=0}^{\inf}\gamma^{k}r_{t+k+1}$ represents the accumulated return from step t. The agent’s goal is to maximize the expected return for each state. The state-action value function $Q^{\pi}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is defined as $Q^{\pi}(s,a)=\mathbb{E}[R_{t}|s_{t}=s,a]$, which is the expected return from selecting action $a$ in state $s_{t}$ following policy $\pi$. The value of state $s$ under policy $\pi$ is defined as $V_{\pi}(s)=\mathbb{E}[R_{t}|s_{t}=s]$, which is the expected return from the state $s$. Policy gradient methods, like REINFORCE Williams (1992) , directly learn a parameterized policy $\pi(s;\theta)$ and update its parameters $\theta$ using approximate gradient ascent of a performance measure, $\nabla_{\theta}J(\theta)=\mathbb{E}_{\pi}[G_{t}\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})]$. To reduce the variance in the expected return estimate, a baseline is subtracted from it; typically, a learned value function is used as the baseline. The resulting quantity $A(s,a)=Q(s,a)-V(s)$ can be viewed as an estimate of the advantage of action $a$ in state $s$. This method is called advantage actor critic (A2C) Sutton and Barto (2018); Degris et al. (2012), and the agent acting according to the policy $\pi$ is the actor, while the value baseline is the critic. Several improved methods have been developed to address sample efficiency, stability, and training difficulty. Proximal policy optimization (PPO) Schulman et al. (2017) is an on-policy policy optimization technique that addresses concerns in actor critic methods using clipped objectives. We adapt PPO for learning policies.
3 Model
@@
Our proposal for metacognitive generalization centers around the ideas of natural decomposition of tasks that an agent can accomplish through interactions with objects. The meta-policy recommends localized agent-object interactions, which the agent executes following interaction policies to construct a compositional solution. Here, we discuss the main ideas of the framework (see Figure 2) and argue that it exhibits task generalization characteristics.
@@
The environment $\Xi$ is partially observable and consists of objects $e=\{e_{1},e_{2},...,e_{n}\}$ and an agent $A$. It is assumed that the agent has a recognition system that can access its partial view $X$ and associate objects with their corresponding types, such that the type of an object $e_{i}\in e$ is in entity types $E=\{E_{1},E_{2},...,E_{m}\}$. The agent can perform low-level actions $u=\{u_{1},u_{2},...\}$ within the environment. When making decisions, the agent organizes its perceived knowledge into knowledge graphs: an instance-level graph and a type-level graph. An instance-level graph is an ordered pair $G_{I}=(e,\Lambda)$ of observed objects $e$ and their relations $\Lambda=\{\Lambda_{1},\Lambda_{2},...,\Lambda_{\kappa}\}$. A type-level graph is an ordered pair $G_{K}=(E,\Lambda)$ of entity types $E$ and their type level or abstract relations $\Lambda$. Relations, whether between objects or entity types, retain the same meaning, such as adjacency indicating that two objects are close to one another. The function $\Phi:X\to G_{I}$ maps the agent’s partially observed view to the instance graph by capturing observed objects and deriving their relationships. The function $\Psi:G_{I}\to G_{K}$ maps instance space graphs to type space graphs, a direct translation since the agent’s recognition system contains types corresponding to the objects. If the agent encounters a novel object without a corresponding type, it must create a new type in its type space or assign the object to a related type. In the framework, $\Psi$ is learnable and represents a belief in mapping objects to types. However, this article treats the mapping as fixed, from object nodes in the instance graph $G_{I}$ to type nodes in the type graph $G_{K}$. As the agent interacts with objects, the environment changes, and its recognition system captures a new partially observed view, resulting in new instance and type space graphs.
The agent’s type space level is also referred to as the meta-level. At the meta-level, the agent uses a meta-policy that accepts a type space knowledge graph and returns meta-level actions or interactions for object interaction. Interactions are represented as $\lambda=\{\lambda_{1},\lambda_{2},...,\lambda_{p}\}$. Similarly, for each interaction type, the agent uses an interaction policy to guide low-level actions $u$ within the environment. Therefore, the agent’s interaction level $\mathcal{I}=(A,G_{I},\lambda)$ comprises of its instance space knowledge graph and available interactions, while the meta-level knowledge $\mathcal{K}=(A,G_{K},\lambda)$ comprises of its type space knowledge graph and interactions.
@@
After the agent constructs the knowledge graphs, the meta-level recommender $\Upsilon$ recommends a candidate object $e_{i}$ and interaction type $\lambda_{p}$ for an interaction based on the agent’s knowledge graphs and meta-level policy $\pi$. The recommender iterates through all visible objects in the agent’s instance graph $G_{I}$. For each object, it adds an activation relation between the agent and the object in both the instance space graph $G_{I}$ and the type space graph $G_{K}$, where ‘activating’ refers to observing or focusing on an object for potential interaction. For $G_{K}$, the agent’s meta-policy returns the state value and interaction probability. The state value represents the value of activating that specific object. Then, an object $e_{i}$ for interaction is proposed through sampling based on the values of activation of the visible objects, and interaction $\lambda_{p}$ is sampled from the interaction probability of the proposed object. Therefore, at each step, the meta-level of the agent activates a candidate object $e_{i}$ for interaction $\lambda_{p}$, setting the agent’s interaction goal $g^{i}_{p}=(e_{i},\lambda_{p})$. Acting on the recommended interaction goal, the agent begins interacting with the environment. After each round of interaction, the interaction status is passed to the meta-level, which then proposes the next candidate object and interaction. This process continues until the task is completed or the environment is terminated.
@@
Each agent has a single meta-level policy, whereas interaction policies are a collection of policies corresponding to different interactions. By following the interaction goals $g^{i}_{p}$ as directed by the meta-level, the agent interacts with objects. During an interaction $\lambda_{p}$, the agent executes a series of low-level actions $u$ on execution level states $s$ based on the relevant interaction policy $\mu_{p}$. For example, if there is a door, the agent can activate it and interact with it, like ‘open’ the door, even if this requires multiple low-level actions such as moving and toggling. As the agent interacts with objects, it collects trajectories of experiences at both the meta and interaction levels. The meta-level trajectory is used to learn the meta-level policy, while the low-level interaction experiences are used to learn the collection of interaction policies. The learning approach for the meta-policy and interaction policies is detailed below.
@@
Meta-level representations of type space graphs with activation relations are learned using a message-passing graph neural network. Nodes have abstract level attributes such as type encoding, while edges have edge type encoding for different types of relationships. We use GAT convolutional layers Brody et al. (2021); Fey and Lenssen (2019) to learn graph representations given as,
$h_{v}^{\prime}=\alpha_{v,v}\theta h_{v}+\sum_{u\in\mathcal{N}(v)}\alpha_{u,v}\theta h_{u}$
@@
$\mathcal{L}_{m}^{\text{actor}}(\theta)=-\mathbb{E}_{\mathcal{D}_{\tau_{m}}}\left[\min\left(r_{\theta}A_{m},\text{clip}(r_{\theta},1-\epsilon,1+\epsilon)A_{m}\right)\right]$
@@
$\mathcal{L}_{m}^{\text{critic}}(\theta)=\mathbb{E}_{\mathcal{D}_{\tau_{m}}}\left[\frac{1}{2}(V_{\pi_{\theta}}(G_{K})-R_{m})^{2}\right]$
@@
$\mathcal{L}_{m}(\theta)=\mathcal{L}_{m}^{\text{actor}}(\theta)-\zeta_{m}^{h}\mathcal{H}_{\theta}+\zeta_{m}^{c}\mathcal{L}_{m}^{\text{critic}}(\theta)$
@@
Interaction policies are a repository of policies for different interactions which can be recommended by the meta-level, $\mu_{\phi}=\{\mu_{\phi_{1}},\mu_{\phi_{2}},...,\mu_{\phi_{p}}\}$ and each policy is parameterized separately. When an interaction goal with an object $e_{i}$ and interaction $\lambda_{p}$ is recommended, $g^{i}_{p}=(e_{i},\lambda_{p})$, the agent attempts to achieve it by executing a series of low-level actions from $u$. The agent observes a low-level state $s$ from a convolutional neural network embedding, samples an action $u\sim\mu_{\phi_{p}}(.|s)$ using the respective interaction policy, acts in the environment, and collects the trajectory of experiences $\mathcal{D}_{\tau_{p}}$. The experiences collected are used to learn interaction policies. An estimate of advantages is computed as $A_{p}=r_{p}+\gamma_{p}V_{\mu_{\phi_{p}}}(s^{\prime})-V_{\mu_{\phi_{p}}}(s)$.
@@
$\mathcal{L}_{p}^{\text{actor}}(\phi_{p})=-\mathbb{E}_{\mathcal{D}_{\tau_{p}}}\left[\min\left(r_{\phi_{p}}A_{p},\text{clip}(r_{\phi_{p}},1-\epsilon,1+\epsilon)A_{p}\right)\right]$

### v3

@@ Abstract
People aptly exhibit general intelligence behaviors through flexible problem-solving and the ability to adapt to novel situations by reusing and applying high-level knowledge acquired over time. In contrast, artificial agents tend to be specialists, lacking such generalist behaviors. To bridge this gap, artificial agents will require understanding and exploiting critical structured knowledge representations. We introduce a metacognitive reasoning framework, Knowledge-Interaction-eXecution (KIX), and argue that interactions with objects, by leveraging a type space, facilitate the learning of transferable interaction concepts and promote generalization. This framework offers a principled approach for integrating knowledge into reinforcement learning and holds promise as an enabler for generalist behaviors in artificial intelligence, robotics, and autonomous systems.
Introduction
@@
The ability to actively reason, flexibly evaluate options and learn from experience is a hallmark of human intelligence. Core to this intelligence is the human ability to identify natural decompositions within complex environments and tasks with ease, enabling flexible and generalist behaviors. Replicating such autonomy and general intelligence abilities in artificial agents continues to pose significant challenges, which we argue necessitate a knowledge and interaction-centric metacognitive approach.
@@
Meta-reasoning about next actions, their values, and possible next states underpins flexible decision-making in novel situations, problem solving, concept learning, and knowledge acquisition. This form of reasoning extends beyond mere memory lookup. More broadly, meta-reasoning functions as a sense-making process within a knowledge system, enabling people to intuitively direct their learning to discern relations between entities, interact with the same entities in diverse ways, and ground these interactions for both short- and long-term goals. Metacognitive reasoning thus emerges as a foundational mechanism for generalization. In contrast, meta-reasoning systems in artificial intelligence are primarily focused on resource allocation problems. While resource management is essential for efficient functioning, it is crucial to ideate metacognitive generalization, illustrated in Figure 1, to enable generalist behaviors in artificial agents.
(a) Overview
(b) Approach
Figure 1: a) Illustration of metacognitive reasoning process in a navigation domain, where an agent (depicted as a red triangle) must interpret its environment to achieve the goal of retrieving a blue ball located behind a locked door and various obstacles. This process requires reasoning over abstract object types and their relations, which in turn guide the agent’s interactions with objects. Such abstract reasoning fundamentally relies on higher-order abstract structures and compositional interactions over these abstractions. b) Learning generalist behaviors requires developing a natural and abstract understanding of the environment, as well as meta-reasoning and interaction abilities. The metacognitive knowledge-guided approach exploits inherent abstractions and relations between objects to generate interaction goals, enabling the agent to interact with objects and generalize to solve complex tasks.
Metacognitive reasoning is fundamentally grounded in interaction, requiring agents to draw upon their interpretations of abstract knowledge to guide their interactions. Interpretation precedes and shapes interaction; it involves developing both natural and abstract understandings of the environment that can be deployed as needed to make sense of the current state and plan subsequent interactions. To illustrate this process, consider the example in Figure 1a, which depicts an agent and its interactions in a navigation domain. To build knowledge of the environment, an agent must develop a conceptual understanding of the objects in the environment and how to interact with them. As some objects may even appear similar, the agent must decipher their abstract types. Learning to interact then involves learning which interactions are possible with these abstract object types and the relations between them. Type serves as a natural abstraction, and interpretation thus incorporates this meta-level reasoning over abstract knowledge. The utility of interactions is then learned through meta-level policies in this abstract space, which can subsequently be leveraged to recommend interactions and execute compositions of interactions at a lower level, thereby facilitating problem-solving.
@@
Our proposal is based on three main principles for flexibly solving complex tasks. Natural abstraction: Real-world environments are naturally composed of entities categorized by their types, similar to a natural hierarchy of objects and classes. Entities in the instance space inherit properties and relations from a higher-level type space. Type-space graphs represent the underlying objects and their relations in environments. This abstraction is where the knowledge resides and it is the knowledge or meta level. Interactions: An agent’s metacognitive process employs the type-space graphs to create a plan specifying which entities to interact with and the nature of these interactions to complete tasks and achieve goals. While the metacognitive policy guides agents based on abstract knowledge, agents interact with entities in the instance space which are instantiations of abstract types. This is the interaction level. Composition of interactions: The agent interacts with meta-guided objects as a recommended interaction type and executes low-level policies corresponding to a guided interaction type. Therefore, a composition of these interactions as interaction pathways for graph traversal at both the meta level and the instance level facilitates execution of low-level actions which are loosely coupled with abstract knowledge through interactions.
@@
Integrating these principles into metacognitive learning, reasoning, and control provides the levers and connecting points for instilling generalist behaviors in agents. Agents’ flexible behaviors stem from their ability to interpret knowledge, learn how to interact with objects through meta level policies and leverage interaction pathways enabling quick adaptation. The proposed metacognitive framework captures the essence of generalist behaviors and may have far-reaching implications across various domains. We describe the key components of the framework that exemplify generalist characteristics, followed by an experimental evaluation in a navigation domain and discuss how we envision it enabling flexibility through meta-level guided interactions.
Background
@@
Compositionality and abstractions are central to generalization learning, with abstraction requiring a mechanism for grounding representations in actual objects. An object-centric model Veerapaneni et al. (2020) was presented in which object representations were constructed from raw observations and simpler local functions with uniformity, which were then used for prediction and planning. Similarly, a symbolic object-centric representation Vijay et al. (2019) was used to study the impact of adding prior knowledge through relations in an object graph. These studies focus primarily on object space, which inhibits the abstraction of objects and relational structures naturally occurring in the real world and used for decision-making. Key questions include what a natural abstraction should be and how it can be used to compose interactions as building blocks leveraging natural relational structures. We address these largely overlooked questions by preserving abstract knowledge and relational higher-order structures in the form of knowledge graphs, where abstractions are grounded using type space. Learning a compositional structure is achieved through learning over knowledge graphs with neighborhood aggregation.
Hierarchical reinforcement learning approaches Dayan and Hinton (1992); Sutton et al. (1999); Pateria et al. (2021); Bagaria et al. (2021) learn a higher-level policy by choosing subtasks as higher-level actions, which decompose tasks into subtasks. Higher-level actions persist over a longer timescale, referred to as temporal abstraction Barto and Mahadevan (2003); Dietterich (2000); Sutton et al. (1999). In contrast to options Sutton et al. (1999), which combine a subtask with the core MDP, MAXQ value function decomposition Dietterich (2000) decomposes the core MDP into smaller sub-MDPs, each with a separate learnable policy. Other methods induce learning hierarchies based on teacher signals Portelas et al. (2020); Hutsebaut-Buysse et al. (2022) or a progression from easier to more difficult tasks Campero et al. (2020). These methods enable agents to learn hierarchy of policies, but they remain in the object space for both low- and high-level actions and do not meet a key meta-reasoning demand: type abstractions and loose coupling between abstraction and execution space. Methods such as Ten et al. (2022); Oudeyer et al. (2007) encourage exploration through intrinsic reward or motivation, Klyubin et al. (2005) develop internal goals for empowerment, focusing on aspects different than ours. On the other hand, neuro-symbolic reasoning approaches Hitzler and Sarker (2022); Zhang and Sheng (2024) aim to overcome the limitations of purely symbolic or purely neural methods by integrating them into hybrid models that combine the learning capacity of neural networks with the logical inference and interpretability of symbolic systems. However, significant challenges persist in facilitating interaction and integration between these disparate paradigms to ensure a seamless, flexible, and semantically grounded natural unification.
Our approach meets these demands by leveraging natural higher-order knowledge structures in the form of type space graphs for meta-level policy and facilitating interactions with objects via an intermediary between meta-level and execution space, thus enabling loose coupling and reuse of low-level policies. The appeal of our approach is that interaction policies can be swapped in and out depending on the interaction type. Our work distinguishes itself by incorporating relational abstract knowledge structures to guide interactions that govern low-level execution.
@@
Meta-reasoning is often defined as the set of processes that monitor our thought processes and regulate resource allocation Ackerman and Thompson (2017). Its basic structure Nelson (1990) consists of two levels: meta and object. The meta level’s functions include control and monitoring. The object’s state changes as control passes from the meta to the object level. The monitoring process allows the meta controller to track the status of object-level processes and actions, as well as the consequences of those actions. However, current meta-reasoning, primarily as a resource manager, is insufficient for exhibiting generalist behaviors, necessitating generalist meta-cognitive approaches such as the one proposed here.
@@
$h_{v}^{(k)}=\mathcal{F}^{\prime}\left(h_{v}^{(k-1)},\bigoplus_{u\in\mathcal{N}(v)}\mathcal{F}^{\prime\prime}\left(h_{v}^{(k-1)},h_{u}^{(k-1)},\mathtt{e}_{u,v}\right)\right)$ (1)
where $h_{v}^{(k)}$ is the updated representation of node $v$ embedding in layer $k$, $\mathcal{N}(v)$ is the set of neighboring nodes of node $v$, $h_{u}^{(k-1)}$ is the representation of neighboring node $u$ in layer $k-1$, $\bigoplus$ is a differentiable permutation invariant aggregation function, $\mathcal{F}^{\prime}$ and $\mathcal{F}^{\prime\prime}$ are differentiable functions usually neural networks. Node features are transformed prior to aggregation, and the methods used to transform and aggregate features differ between graph network variants. Motivated by the attention mechanism Vaswani et al. (2017) in the transformer model, Veličković et al. (2017) employs self-attention over node features and aggregating over neighborhood nodes in graph attention networks (GAT). We use the GATv2 operator Brody et al. (2021), which improves static attention problems in graph attention layers.
@@
We consider a standard reinforcement learning setting Sutton and Barto (2018) in which an agent interacts with the environment over a number of steps. A Markov decision process (MDP) is defined as a tuple $(\mathcal{S},\mathcal{A},\mathcal{P},r,\gamma)$ in which $\mathcal{S}$ is a finite state space, $\mathcal{A}$ is a finite discrete action space, $\mathcal{P}:\mathcal{S}\times\mathcal{A}\times\mathcal{S}\rightarrow\mathbb{R}$ is the transition model, $r:\mathcal{S}\times\mathcal{S}\rightarrow\mathbb{R}$ is the reward function and $\gamma\in(0,1)$ is the discount factor. At each step t, the agent takes an action $a_{t}\in\mathcal{A}$ in state $s_{t}\in\mathcal{S}$ according to its policy $\pi$, which is a mapping of states to actions. As a result, the agent reaches the state $s_{t+1}\in\mathcal{S}$, receives a scalar reward $r_{t}\in\mathbb{R}$, and accumulates returns; $R_{t}=\sum_{k=0}^{\inf}\gamma^{k}r_{t+k+1}$ represents the accumulated return from step t. The agent’s goal is to maximize the expected return for each state. The state-action value function $Q^{\pi}:\mathcal{S}\times\mathcal{A}\rightarrow\mathbb{R}$ is defined as $Q^{\pi}(s,a)=\mathbb{E}[R_{t}|s_{t}=s,a]$, which is the expected return from selecting action $a$ in state $s_{t}$ following policy $\pi$. The value of state $s$ under policy $\pi$ is defined as $V_{\pi}(s)=\mathbb{E}[R_{t}|s_{t}=s]$, which is the expected return from the state $s$. Policy gradient methods, like REINFORCE Williams (1992), directly learn a parameterized policy $\pi(s;\theta)$ and update its parameters $\theta$ using approximate gradient ascent of a performance measure, $\nabla_{\theta}J(\theta)=\mathbb{E}_{\pi}[G_{t}\nabla_{\theta}\ln\pi_{\theta}(A_{t}|S_{t})]$. To reduce the variance in the expected return estimate, a baseline is subtracted from it; typically, a learned value function is used as the baseline. The resulting quantity $A(s,a)=Q(s,a)-V(s)$ can be viewed as an estimate of the advantage of action $a$ in state $s$. This method is called advantage actor critic (A2C) Sutton and Barto (2018); Degris et al. (2012), and the agent acting according to the policy $\pi$ is the actor, while the value baseline is the critic. Several improved methods have been developed to address sample efficiency, stability, and training difficulty. Proximal policy optimization (PPO) Schulman et al. (2017) is an on-policy policy optimization technique that addresses concerns in actor-critic methods using clipped objectives. We adapt PPO for learning policies.
Metacognitive framework
@@
Our proposal for metacognitive generalization centers around the ideas of natural decomposition of tasks that an agent can accomplish through interpretation of knowledge and interpretation guided interactions with objects. The meta-policy recommends localized agent-object interactions, which the agent executes following interaction policies to construct a compositional solution. Here, we discuss the main ideas of the framework (see Figure 2) and argue that it exhibits task-generalization characteristics.
@@
The environment $\Xi$ is partially observable and consists of objects $e=\{e_{1},e_{2},...,e_{n}\}$ and an agent $A$. It is assumed that the agent has a recognition system that can access its partial view $X$ and associate objects with their corresponding types, such that the type of an object $e_{i}\in e$ is in entity types $E=\{E_{1},E_{2},...,E_{m}\}$. The agent can perform low-level actions $u=\{u_{1},u_{2},...\}$ within the environment. When making decisions, the agent organizes its perceived knowledge into graphs: an instance-level graph and a type-level graph. An instance-level graph is an ordered pair $G_{I}=(e,\Lambda)$ of observed objects $e$ and their relations $\Lambda=\{\Lambda_{1},\Lambda_{2},...,\Lambda_{\kappa}\}$. A type-level graph is an ordered pair $G_{K}=(E,\Lambda)$ of entity types $E$ and their type level or abstract relations $\Lambda$. Relations, whether between objects or entity types, retain the same meaning, such as adjacency indicating that two objects are close to one another. The function $\Phi:X\to G_{I}$ maps the agent’s partially observed view to the instance graph by capturing observed objects and deriving their relationships. The function $\Psi:G_{I}\to G_{K}$ maps instance space graphs to type space graphs, a direct translation since the agent’s recognition system contains types corresponding to the objects. If the agent encounters a novel object without a corresponding type, it must create a new type in its type space or assign the object to a related type. In the framework, $\Psi$ is learnable and represents a belief in mapping objects to types. However, this article treats the mapping as fixed, from object nodes in the instance graph $G_{I}$ to type nodes in the type graph $G_{K}$. As the agent interacts with objects, the environment changes, and its recognition system captures a new partially observed view, resulting in new instance and type space graphs.
The agent’s knowledge level is also referred to as the meta-level. At the meta-level, the agent uses a meta-policy that accepts a type space graph and returns interactions or meta-level actions for object interaction. Interactions are represented as $\lambda=\{\lambda_{1},\lambda_{2},...,\lambda_{p}\}$. Similarly, for each interaction type, the agent uses an interaction policy to guide low-level actions $u$ within the environment. Therefore, the agent’s interaction level $\mathcal{I}=(A,G_{I},\lambda)$ comprises of its instance space graph and available interactions, while the meta or knowledge level $\mathcal{K}=(A,G_{K},\lambda)$ comprises of its type space graph and meta-level actions or interactions.
@@
After the agent constructs the type and instance graphs, the meta-level recommender $\Upsilon$ recommends a candidate object $e_{i}$ and interaction type $\lambda_{p}$ for an interaction based on the agent’s type and instance graphs and meta-level policy $\pi$. The recommender iterates through all visible objects in the agent’s instance graph $G_{I}$. For each object, it adds an activation relation between the agent and the object in both the instance space graph $G_{I}$ and the type space graph $G_{K}$, where ‘activating’ refers to observing or focusing on an object for potential interaction. For $G_{K}$, the agent’s meta-policy returns the state value and the interaction probability. The state value represents the value of activating that specific object. Then, an object $e_{i}$ for interaction is proposed by sampling based on the values of activation of the visible objects, and interaction $\lambda_{p}$ is sampled from the interaction probability of the proposed object. Therefore, at each step, the meta-level of the agent activates a candidate object $e_{i}$ for interaction $\lambda_{p}$, setting the agent’s interaction goal $g^{i}_{p}=(e_{i},\lambda_{p})$. Acting on the recommended interaction goal, the agent begins interacting with the environment. After each round of interaction, the interaction status is passed to the meta-level, which then proposes the next candidate object and interaction. This process continues until the task is completed or the environment is terminated.
@@
Each agent has a meta-level policy, whereas interaction policies are a collection of policies corresponding to different interactions. By following the interaction goals $g^{i}_{p}$ as directed by the meta-level, the agent interacts with objects. During an interaction $\lambda_{p}$, the agent executes a series of low-level actions $u$ on execution level states $s$ based on the relevant interaction policy $\mu_{p}$. For example, if there is a door, the agent can activate it and interact with it, like ‘open’ the door, even if this requires multiple low-level actions such as moving and toggling. As the agent interacts with objects, it collects trajectories of experiences at both the meta and interaction levels. The meta-level trajectory is used to learn the meta-level policy, while the low-level interaction experiences are used to learn the collection of interaction policies. The learning approach for the meta-policy and interaction policies is detailed below.
@@
Meta-level representations of type space graphs with activation relations are learned using a message-passing graph neural network. Nodes have abstract level attributes such as type encoding, while edges have edge-type encoding for different types of relations. We use GAT convolutional layers Brody et al. (2021); Fey and Lenssen (2019) to learn graph representations given as,
$h_{v}^{\prime}=\alpha_{v,v}\theta h_{v}+\sum_{u\in\mathcal{N}(v)}\alpha_{u,v}\theta h_{u}$ (2)
@@
$\mathcal{L}_{m}^{\text{actor}}(\theta)=-\mathbb{E}_{\mathcal{D}_{\tau_{m}}}\left[\min\left(r_{\theta}A_{m},\text{clip}(r_{\theta},1-\epsilon,1+\epsilon)A_{m}\right)\right]$ (3)
@@
$\mathcal{L}_{m}^{\text{critic}}(\theta)=\mathbb{E}_{\mathcal{D}_{\tau_{m}}}\left[\frac{1}{2}(V_{\pi_{\theta}}(G_{K})-R_{m})^{2}\right]$ (4)
@@
$\mathcal{L}_{m}(\theta)=\mathcal{L}_{m}^{\text{actor}}(\theta)-\zeta_{m}^{h}\mathcal{H}_{\theta}+\zeta_{m}^{c}\mathcal{L}_{m}^{\text{critic}}(\theta)$ (5)
@@
Interaction policies are a repository of policies for different interactions which can be recommended by the meta level, $\mu_{\phi}=\{\mu_{\phi_{1}},\mu_{\phi_{2}},...,\mu_{\phi_{p}}\}$ and each policy is parameterized separately. When an interaction goal with an object $e_{i}$ and interaction $\lambda_{p}$ is recommended, $g^{i}_{p}=(e_{i},\lambda_{p})$, the agent attempts to achieve it by executing a series of low-level actions from $u$. The agent observes a low-level state $s$ from a convolutional neural network embedding, samples an action $u\sim\mu_{\phi_{p}}(.|s)$ using the respective interaction policy, acts in the environment, and collects the trajectory of experiences $\mathcal{D}_{\tau_{p}}$. The experiences collected are used to learn interaction policies. An estimate of advantages is computed as $A_{p}=r_{p}+\gamma_{p}V_{\mu_{\phi_{p}}}(s^{\prime})-V_{\mu_{\phi_{p}}}(s)$.
@@
$\mathcal{L}_{p}^{\text{actor}}(\phi_{p})=-\mathbb{E}_{\mathcal{D}_{\tau_{p}}}\left[\min\left(r_{\phi_{p}}A_{p},\text{clip}(r_{\phi_{p}},1-\epsilon,1+\epsilon)A_{p}\right)\right]$ (6)